OpenStudy (anonymous):

Is 3+(-7)=10 a true false or open sentence. I am trying to help my 7th grader with his math

OpenStudy (anonymous):

false it would equal -4

OpenStudy (anonymous):

3 + (-7) = 10 3 - 7 = 10 -4 = 10 -------> FALSE :)

OpenStudy (mathteacher1729):

@jordyn77 the left side would equal -4. 3 + (-7) = -4 But the rights side is 10 and -4 does not equal 10, so the statement is false, as @kmeds16 said. :)

OpenStudy (anonymous):

Well, um I said that.. ? It's the same concept.

OpenStudy (anonymous):

-4 is not equal to 10